# Forget Vectors at Play: Universal Input Perturbations Driving Machine Unlearning in Image Classification

### v2

@@ 4 Generalization of MU to Forget Data Shifts
(a) Data Corruptions. Following the OOD generalization evaluation approach in image classification (Hendrycks et al., 2021; Hendrycks & Dietterich, 2019), we introduce data corruptions from four main categories: noise, blur, weather, and digital. Each type of corruption includes five levels of severity, with higher levels representing increased noise intensity. Among these, we select zero-mean Gaussian noise (GN) and Elastic transformations (ET) as the primary corruption types to evaluate MU robustness against shifts in forget data. Our rationale is that Gaussian noise yields small pixel-wise perturbations (similar to adversarial perturbations introduced later) and Elastic transformations stretch or contract small image regions.
@@
As shown in Fig. 2-(a) and (b), model utility, measured by RA (retain accuracy) and TA (testing accuracy), decreases when external perturbations are applied to the evaluation sets compared to its original performance without perturbations. This is expected due to the generalization loss when facing data shifts. More interestingly, Fig. 2-(c) and (d) show that unlearning effectiveness of Retrain, measured by UA (unlearning accuracy) and MIA-Efficacy, remains stable despite the presence of these perturbations on the forget set. This is because perturbations degrade prediction performance across evaluation sets, including the forget set. This is further evidenced by the increase in UA and MIA-Efficacy for the original model (without unlearning) when exposed to data perturbations. The above indicates that a reduction in prediction performance on the forget set could translate into enhanced unlearning effectiveness on that set. In Appendix A, we provide additional evaluations of other approximate unlearning methods, including FT, RL, and GA, showing consistent performance.
The results above demonstrate that unlearning effectiveness is inherently preserved under external perturbations at no additional cost. However, balancing this with utility retention in the presence of perturbations remains challenging. Therefore, we need to address the forget vector problem (4) to develop an input-based MU solution that enhances unlearning effectiveness without compromising model utility.
@@ 5 Optimization for Forget Vectors
As motivated in Sec. 4, if we treat the forget vector as an external perturbation applied to the forget set, our goal becomes optimizing its design to maximize unlearning effectiveness without compromising the original model utility. To this end, we first propose an unlearning objective function, $\ell_{\mathrm{MU}}$, tailored for the forget vector design in (4), inspired by the untargeted C&W attack (Carlini & Wagner, 2017). We then introduce a novel paradigm called compositional unlearning, facilitated by forget vector arithmetic. These approaches center on optimizing the input perturbation $\bm{\delta}$ in (4) while keeping the model unchanged, i.e., identical to the pre-trained original model ${\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}$.
@@
Our design aims for the forget vector variable ($\bm{\delta}$), when applied to the forget set ($\mathcal{D}_{\mathrm{f}}$), to drive the model’s predictions (${{{\bm{\theta}}_{\mathrm{o}}}}$) away from the correct labels. Conversely, when applied to the retain set ($\mathcal{D}\mathrm{r}$), the forget vector should minimally affect correct predictions. The first forget objective aligns with adversarial attack design, aiming to mislead the model’s predictions in the presence of the perturbation $\bm{\delta}$. The second retain objective acts as a utility regularization, suppressing the unlearning effect of the perturbation when applied to data samples not targeted for unlearning.
To implement the forget objective (denoted by $\ell_{\mathrm{f}}$), we draw inspiration from the C&W untargeted attack loss (Carlini & Wagner, 2017). This is given by a margin loss, designed to remain active when the top prediction matches the correct label, ensuring that optimization continues until predictions are shifted to an incorrect label, thereby achieving unlearning. This yields
@@
where $(\mathbf{x},y)\in\mathcal{D_{\mathrm{f}}}$ denotes a forget sample with $y$ being the true label of $\mathbf{x}$, $\mathbf{x}+\bm{\delta}$ is the perturbed sample, $f_{k,{{\bm{\theta}}_{\mathrm{o}}}}(\mathbf{x})$ denotes the prediction logit (before softmax) of the model ${{\bm{\theta}}_{\mathrm{o}}}$ for class $k$ under the input $\mathbf{x}$, and $\tau\geq 0$ is a margin threshold that controls the unlearning strength. The rationale behind (5) is that minimizing it ensures convergence to the negative margin $f_{{y}}(\mathbf{x}+\bm{\delta})-\max_{{k}\neq{y}}f_{k}(\mathbf{x}+\bm{\delta})\to-\tau\leq 0$. Thus, the forget vector $\bm{\delta}$ enforces unlearning on $\bm{\theta}$ for $\mathbf{x}$ by making the incorrect prediction ($k\neq y$) have a higher confidence than the original correct prediction $y$. On the other hand, once the margin becomes negative (indicating that the prediction label has been flipped), the forget objective $\ell_{\mathrm{f}}$ automatically terminates, allowing a balance with the retain objective, which will be introduced later. In our experiments, we find that the forget objective is robust to variations in the nonnegative margin parameter $\tau$ (see Appendix B). A larger $\tau$ value imposes a stricter unlearning requirement by increasing the logit distance from the correct label. For example, we set $\tau=1$ in our experiments.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) UA1 (b) UA2 (c) TA (d) Avg. Performance
Figure 3: The performance of forget vector arithmetic regarding class 1 and class 2 on CIFAR-10 with ResNet-18 evaluated across different combination coefficients $w_{1}$ and $w_{2}$ ranging from 00 to $1$ with an interval of $0.2$, includes (1) unlearning accuracy for class 1 (UA1), (2) unlearning accuracy for class 2 (UA2), (3) testing accuracy for remaining classes (TA), and (4) the average (avg.) performance across UA1, UA2, and TA, with a green star () marking the optimized weight scheme ($w_{1}$ and $w_{2}$) that minimizes (6) over $\mathbf{w}$ in (7). The color bar on the right represents a gradient scale from light to dark red, indicating the range of values (00 to $100$%) in the heatmap. The integer within each cell represents the performance (%) given a combination of $w_{1}$ and $w_{2}$.
@@
A forget vector defines an unlearning direction in the input space to guide the unlearning process. We explore whether a new unlearning direction can be efficiently constructed by interpolating from existing forget vectors, such as class-wise forget vectors obtained by solving (6) with $\mathcal{D}_{\mathrm{f}}$ defined as each class’s training set. This approach is analogous to the concept of task vectors in weight space for model editing (Ilharco et al., 2022). However, to the best of our knowledge, input-based task vector arithmetic has not yet been explored in the literature. If forget vectors can be modified and combined using arithmetic operations, such as negation and addition, we can dynamically adjust a model’s unlearning behavior without re-solving the optimization problem (6) or any other model-based unlearning problem in (2). We refer to this new unlearning paradigm as compositional unlearning.
@@
To illustrate the effectiveness of forget vector arithmetic, Fig. 3 shows preliminary results of combining two class-wise forget vectors ($\bm{\delta}_{1}$ and $\bm{\delta}_{2}$) using a simple scheme $\bm{\delta}(\mathbf{w})=w_{1}\bm{\delta}_{1}+w_{2}\bm{\delta}_{2}$ on (CIFAR-10, ResNet-18). Rather than optimizing $\mathbf{w}$, we adjust its value and observe changes in unlearning accuracy for image class 1 (UA1) and for class 2 (UA2), as well as TA (testing accuracy) for the remaining classes, as $w$ varies. As shown in Fig. 3-(a) and (b), higher UA1 and UA2 values occur with increased $w_{1}$ or $w_{2}$, indicating a positive impact on unlearning performance for classes 1 and 2. Additionally, Fig. 3-(c) reveals that TA remains relatively stable under moderate weighting. These results suggest that more aggressive weighting of forget vectors enhances unlearning for the target classes but may negatively impact testing accuracy for other classes. An optimal weighting scheme can be identified by minimizing problem (6) over $\mathbf{w}$ using (7). This is supported by the average performance shown in Fig. 3-(d), where the highest average performance aligns closely with the optimized weighting scheme marked by the green star ().
@@ 6.1 Experiment Setups
We focus on MU for image classification, using two datasets: CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-10, a 10-class subset of the original ImageNet (Deng et al., 2009), for ease of implementation of Retrain (exact unlearning) over ImageNet images as (Huang et al., 2022; Tao et al., 2021). For these tasks, we use two well-trained image classifiers: ResNet-18 (He et al., 2016) for CIFAR-10 and VGG-16 (Simonyan & Zisserman, 2014) for ImageNet-10 classification.
@@
In the context of MU for image classification, we consider two scenarios: class-wise forgetting and data-wise forgetting. In class-wise forgetting, training data from an image class are designated for unlearning, while in random data forgetting, a subset of all-class training points is randomly selected as the forget set, with a specified forget ratio of 10%. To demonstrate the effectiveness of our proposal, we consider 6 MU baseline methods, including ① FT (Warnecke et al., 2021), ② RL (Golatkar et al., 2020), ③ GA (Thudi et al., 2022a), ④ NegGrad+ (Kurmanji et al., 2024a), ⑤ SalUn (Fan et al., 2023), and ⑥ SCRUB (Kurmanji et al., 2024b).
As described in Sec. 3, unlearning effectiveness is measured using UA (unlearning accuracy) and MIA-Efficacy, while model utility post-unlearning is assessed by RA (retain accuracy) and TA (testing accuracy); for all metrics, higher values indicate better performance. It is also worth noting that all existing model-based MU baseline methods ①-⑥ are evaluated on non-perturbed evaluation sets. However, when using our proposed data-based forget vector solution, we need to perturb the evaluation sets (including the forget set, retain set, and testing set) through the forget vector to assess unlearning effectiveness and utility retention.
@@
To solve the forget vector problem (4) with the proposed unlearning objective in (6), we set the retain loss regularization parameter $\lambda_{1}$ to 3 for CIFAR-10 and 5 for ImageNet-10 in class-wise forgetting, and to 1 for random data forgetting. The $\ell_{2}$-norm regularization parameter is set to $\lambda_{2}=1$. These hyperparameters are determined through a grid search over the range $[0,10]$. To optimize (4), we use stochastic gradient descent (SGD) (Amari, 1993) with a momentum factor of $0.9$ and an exponential learning rate scheduler, decaying at a rate of $0.9$ per iteration. Additionally, the batch size is set to $256$, with a maximum of 40 optimization iterations for both two datasets. To solve the compositional unlearning problem (7), we use a similar setup, setting both $\lambda_{1}$ and $\lambda_{2}$ to $1$.

### v4

@@ 4 Generalization of MU to Forget Data Shifts
(a) Data Corruptions. Following the OOD generalization evaluation approach in image classification (Hendrycks et al., 2021; Hendrycks & Dietterich, 2019), we consider data corruptions from four main categories: noise, blur, weather, and digital. Each type of corruption includes five levels of severity, with higher levels representing increased noise intensity. Among these, we select zero-mean Gaussian noise (GN) and Elastic transformations (ET) as the primary corruption types to evaluate MU robustness against shifts in forget data. Our rationale is that Gaussian noise yields small pixel-wise perturbations (similar to adversarial perturbations introduced later) and Elastic transformations stretch or contract small image regions.
@@
As shown in fig. 2-(a) and (b), model utility, measured by RA (retain accuracy) and TA (testing accuracy), decreases when external perturbations are applied to the evaluation sets compared to its original performance without perturbations. This is expected due to the generalization loss when evaluated on new, shifted data. More interestingly, fig. 2-(c) and (d) show that unlearning effectiveness of Retrain, measured by UA (unlearning accuracy) and MIA-Efficacy, remains stable despite the presence of these perturbations on the forget set. This is because perturbations degrade prediction performance across evaluation sets, including the forget set. This is further evidenced by the increase in UA and MIA-Efficacy for the original model (without unlearning) when exposed to data perturbations. The above indicates that a reduction in prediction performance on the forget set could translate into enhanced unlearning effectiveness on that set. In Appendix B, we provide additional evaluations of other approximate unlearning methods, including FT, RL, and GA, showing consistent performance.
The results above demonstrate that unlearning effectiveness is inherently preserved under external perturbations at no additional cost. However, balancing this with utility retention in the presence of perturbations remains challenging and desirable. Therefore, we need to carefully address the forget vector problem (4) to develop an input-based MU solution that enhances unlearning effectiveness without compromising model utility.
@@ 5 Optimization for Forget Vectors
In this section, we first propose an unlearning objective function, $\ell_{\mathrm{MU}}$, tailored for the forget vector design in (4), inspired by the untargeted C&W attack (Carlini & Wagner, 2017). We then introduce a novel paradigm called compositional unlearning, facilitated by forget vector arithmetic.
@@
Our design aims for the forget vector variable ($\bm{\delta}$), when applied to the forget set ($\mathcal{D_{\mathrm{f}}}$), to drive the given model’s predictions (${{{\bm{\theta}}_{\mathrm{o}}}}$) away from the correct labels. Conversely, when applied to the retain set ($\mathcal{D_{\mathrm{r}}}$), the forget vector should minimally affect correct predictions. The first forget objective aligns with adversarial attack design, aiming to mislead the model’s predictions in the presence of the perturbation $\bm{\delta}$. The second retain objective acts as a utility regularization, suppressing the unlearning effect of the perturbation when applied to data samples not targeted for unlearning.
To implement the forget objective (denoted by $\ell_{\mathrm{f}}$), we draw inspiration from the C&W untargeted attack loss (Carlini & Wagner, 2017). This is given by a margin loss, designed to remain actively minimizing when the top prediction matches the correct label, ensuring that optimization continues until predictions are shifted to an incorrect label, thereby achieving unlearning. This can be cast as
@@
where $(\mathbf{x},y)\in\mathcal{D_{\mathrm{f}}}$ denotes a forget sample with $y$ being the prediction label of $\mathbf{x}$, $\mathbf{x}+\bm{\delta}$ is the perturbed sample, $f_{k,{{\bm{\theta}}_{\mathrm{o}}}}(\mathbf{x})$ denotes the prediction logit (before softmax) of the model ${{\bm{\theta}}_{\mathrm{o}}}$ for class $k$ under the input $\mathbf{x}$, and $\tau\geq 0$ is a margin threshold that controls the unlearning strength. The rationale behind (5) is that minimizing it ensures convergence to the negative margin $f_{{y}}(\mathbf{x}+\bm{\delta})-\max_{{k}\neq{y}}f_{k}(\mathbf{x}+\bm{\delta})\to-\tau\leq 0$. Thus, the forget vector $\bm{\delta}$ enforces unlearning on $\bm{\theta}$ for $\mathbf{x}$ by making the incorrect prediction ($k\neq y$) have a higher confidence than the original correct prediction $y$. On the other hand, once the margin becomes negative (indicating that the prediction label has been flipped), the forget objective $\ell_{\mathrm{f}}$ automatically terminates, allowing a balance with the retain objective, which will be introduced later. In our experiments, we find that the forget objective is robust to variations in the nonnegative margin parameter $\tau$ (see Appendix C). A larger $\tau$ value imposes a stricter unlearning requirement by increasing the logit distance from the correct label. For example, we set $\tau=1$ in our experiments.
@@
A forget vector defines an unlearning direction in the input space to guide the unlearning process. We explore whether a new unlearning direction can be efficiently constructed by interpolating from existing precomputed forget vectors, such as class-wise forget vectors obtained by solving (6) with $\mathcal{D}_{\mathrm{f}}$ defined as each class’s training set. This approach is analogous to the concept of task vectors in weight space for model editing (Ilharco et al., 2022). However, to the best of our knowledge, input-based task vector arithmetic has not yet been explored in the literature. If forget vectors can be modified and combined using arithmetic operations, such as negation and addition, we can dynamically adjust a model’s unlearning behavior without re-solving the optimization problem (6) or any other model-based unlearning problem in (2). We refer to this new unlearning paradigm as compositional unlearning, where precomputed class-wise forget vectors can be efficiently combined to generate a new forget vector for each deletion request in the context of random data forgetting.
@@
Refer to caption Refer to caption Refer to caption
(a) UA Gap (b) RA Gap (c) Avg. Gap
Figure 3: The performance gap relative to Retrain for class-wise forget vector arithmetic (based on classes “automobile” and “bird”) across different combination coefficients $w_{1}$ and $w_{2}$, when unlearning a randomly selected 10% of training points from these two classes of CIFAR-10. Each cell displays the gap (%) relative to Retrain at a specific weight combination, where a lower value indicates a closer performance to Retrain given a metric. A green star () denotes the selected weight combination scheme ($w_{1}$ and $w_{2}$) that achieves the smallest performance gap relative to Retrain, averaged over both UA Gap and RA Gap.
To illustrate the effectiveness of forget vector arithmetic, Fig. 3 shows preliminary results of combining two class-wise forget vectors ($\bm{\delta}_{1}$ and $\bm{\delta}_{2}$) using a simple scheme $\bm{\delta}(\mathbf{w})=w_{1}\bm{\delta}_{1}+w_{2}\bm{\delta}_{2}$ on (CIFAR-10, ResNet-18), when forgetting a randomly selected 10% of training points from class “automobile” and class “bird” refering to $w_{1}$ and $w_{2}$. Rather than optimizing $\mathbf{w}$, we adjust $w_{1}$ and $w_{2}$ from $-0.2$ to $0.2$ to observe how the performance gap relative to Retrain varies. This evaluation includes the UA (unlearning accuracy) gap on the selected forget data, the RA (retain accuracy) gap on the remaining data, and the average gap across these two metrics. As expected, Fig. 3-(a) and (b) shows a trade-off among these two metrics, where weight configurations that achieve a low UA gap may result in a higher RA gap, and vice versa. Additionally, Fig. 3-(c) shows that moderate weight values of $w_{1}$ and $w_{2}$ ($-0.1\leq w_{1}\leq 0.1$ and $-0.1\leq w_{2}\leq 0.1$ ) tent to yield a more balanced average performance, maintaining relatively low gaps across two metrics. A favored weighting scheme can be identified at $w_{1}=-0.1$ and $w_{2}=0.1$ as marked by the green star (), validating the feasibility of arbitrary random data forgetting using our proposed compositional unlearning via the forget vector arithmetic approach.
@@ 6.1 Experiment Setups
We focus on MU for image classification, using two datasets: CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-10, a 10-class subset of the original ImageNet (Deng et al., 2009), for ease of implementation of Retrain (exact unlearning) over ImageNet images as (Huang et al., 2022; Tao et al., 2021). For these tasks, we use three well-trained image classifiers: ResNet-18 (He et al., 2016) for CIFAR-10, VGG-16 (Simonyan & Zisserman, 2014) and ViT-(Dosovitskiy et al., 2020) for ImageNet classification.
@@
In the context of MU for image classification, we consider two scenarios: class-wise forgetting and data-wise forgetting. In class-wise forgetting, training data from an image class are designated for unlearning, while in random data forgetting, a subset of all-class training points is randomly selected as the forget set, with a specified forget ratio of 10%. To demonstrate the effectiveness of our proposal, we consider $8$ MU baseline methods, including ① FT (Warnecke et al., 2021), ② RL (Golatkar et al., 2020), ③ GA (Thudi et al., 2022a), ④ NegGrad+ (Kurmanji et al., 2024a), ⑤ SalUn (Fan et al., 2023), ⑥ SCRUB (Kurmanji et al., 2024b), ⑦ Class-F (Kodge et al., 2024), and ⑧ SSD (Foster et al., 2024a), where Class-F is only designed for class-wise forgetting.
As described in Sec. 3, unlearning effectiveness is measured using UA (unlearning accuracy) and MIA-Efficacy, while model utility post-unlearning is assessed by RA (retain accuracy) and TA (testing accuracy); For all metrics, being closer to Retrain indicates better performance. It is also worth noting that all existing model-based MU baseline methods ①-⑧ are evaluated on non-perturbed evaluation sets. However, when using our proposed data-based forget vector solution, we need to apply the forget vector to the evaluation sets (including the forget set, retain set, and testing set) in order to assess unlearning effectiveness and utility retention. This evaluation remains fair, as it aligns with the same objective of forgetting targeted data. The key distinction is that the forget vector operates at the input level, whereas model-based MU baselines achieve unlearning by modifying model weights.
@@
To solve the forget vector problem (4) with the proposed unlearning objective in (6), we set the retain loss regularization parameter $\lambda_{1}$ as follows: $3$ for CIFAR-10, $5$ for ImageNet-10 with VGG-16, and $7$ for ImageNet-10 with ViT in class-wise forgetting. For random data forgetting, we set $\lambda_{1}$ to $1$. The $\ell_{2}$-norm regularization parameter is set to $\lambda_{2}=1$. These hyperparameters are determined through a grid search over the range $[0,10]$. To optimize (4), we use stochastic gradient descent (SGD) (Amari, 1993) with a momentum factor of $0.9$ and an exponential learning rate scheduler, decaying at a rate of $0.9$ per iteration. Additionally, the batch size is set to $256$, with a maximum of 40 optimization iterations for both two datasets. To solve the compositional unlearning problem (7), we use a similar setup, setting both $\lambda_{1}$ and $\lambda_{2}$ to $1$.
[…]
Class-wise Forgetting, CIFAR-10, ResNet-18 Random Data Forgetting, CIFAR-10, ResNet-18
Retrain $100.00_{\pm{0.00}}$($0.00$) $100.00_{\pm{0.00}}$($0.00$) $99.91_{\pm{0.03}}$($0.00$) $94.92_{\pm{0.15}}$($0.00$) $\mathbf{0.00}$ $5.50_{\pm{0.16}}$($0.00$) $11.57_{\pm{0.47}}$($0.00$) $99.88_{\pm{0.05}}$($0.00$) $94.24_{\pm{0.19}}$($0.00$) $\mathbf{0.00}$
FT $5.27_{\pm{0.73}}$($94.73$) $51.49_{\pm{5.07}}$($48.51$) $100.0_{\pm{0.0}}$($0.09$) $95.03_{\pm{0.07}}$($0.11$) $\mathbf{35.86}$ $0.03_{\pm{0.03}}$($5.47$) $0.75_{\pm{0.09}}$($10.82$) $99.98_{\pm{0.02}}$($0.10$) $94.45_{\pm{0.14}}$($0.21$) $\mathbf{4.15}$
RL $18.87_{\pm{7.34}}$($81.13$) $98.94_{\pm{0.79}}$($1.06$) $99.98_{\pm{0.0}}$($0.07$) $94.51_{\pm{0.12}}$($0.41$) $\mathbf{20.67}$ $0.52_{\pm{0.24}}$($4.98$) $3.13_{\pm{0.55}}$($8.44$) $99.85_{\pm{0.07}}$($0.03$) $93.88_{\pm{0.20}}$($0.36$) $\mathbf{3.45}$
GA $71.45_{\pm{0.35}}$($28.55$) $81.7_{\pm{0.22}}$($18.30$) $98.62_{\pm{0.04}}$($1.29$) $92.34_{\pm{0.02}}$($2.58$) $\mathbf{12.68}$ $1.56_{\pm{3.08}}$($3.94$) $2.88_{\pm{3.44}}$($8.69$) $98.67_{\pm{2.74}}$($1.21$) $92.84_{\pm{2.59}}$($1.40$) $\mathbf{3.81}$
NegGrad+ $91.78_{\pm{14.66}}$($8.22$) $95.81_{\pm{7.28}}$($4.19$) $98.35_{\pm{1.22}}$($1.56$) $92.62_{\pm{1.34}}$($2.30$) $\mathbf{4.07}$ $0.97_{\pm{1.08}}$($4.53$) $2.74_{\pm{2.16}}$($8.83$) $99.42_{\pm{0.87}}$($0.46$) $93.38_{\pm{1.13}}$($0.86$) $\mathbf{3.67}$
SalUn $96.35_{\pm{2.14}}$($3.65$) $98.64_{\pm{0.03}}$($1.36$) $98.75_{\pm{0.18}}$($1.16$) $92.34_{\pm{1.54}}$($2.58$) $\mathbf{2.19}$ $1.73_{\pm{0.25}}$($3.77$) $6.25_{\pm{1.21}}$($5.32$) $99.24_{\pm{0.09}}$($0.64$) $91.03_{\pm{0.14}}$($3.21$) $\mathbf{3.23}$
SCRUB $93.45_{\pm{2.33}}$($6.55$) $96.38_{\pm{1.72}}$($3.62$) $99.95_{\pm{0.0}}$($0.04$) $94.56_{\pm{0.07}}$($0.36$) $\mathbf{2.64}$ $0.61_{\pm{0.31}}$($4.89$) $3.69_{\pm{0.54}}$($7.88$) $99.76_{\pm{0.18}}$($0.12$) $93.91_{\pm{0.19}}$($0.33$) $\mathbf{3.31}$
Class-F $90.18_{\pm{1.20}}$($9.82$) $86.15_{\pm{2.67}}$($13.85$) $91.25_{\pm{0.05}}$($8.66$) $85.45_{\pm{0.19}}$($9.47$) $\mathbf{10.45}$ n/a n/a n/a n/a n/a
SSD $96.05_{\pm{0.45}}$($3.95$) $98.00_{\pm{0.00}}$($2.00$) $97.77_{\pm{0.20}}$($2.14$) $92.23_{\pm{0.58}}$($2.69$) $\mathbf{2.70}$ $5.54_{\pm{0.00}}$($0.04$) $7.80_{\pm{0.00}}$($3.77$) $94.86_{\pm{0.00}}$($5.02$) $88.28_{\pm{0.00}}$($5.96$) $\mathbf{3.70}$
Ours $97.88_{\pm{0.27}}$($2.12$) $99.60_{\pm{0.15}}$($0.40$) $97.25_{\pm{0.24}}$($2.66$) $90.90_{\pm{0.32}}$($4.02$) $\mathbf{2.30}$ $2.61_{\pm{0.49}}$($2.89$) $8.26_{\pm{1.17}}$($3.00$) $97.33_{\pm{0.47}}$($2.55$) $90.97_{\pm{0.38}}$($3.27$) $\mathbf{2.92}$
Class-wise Forgetting, ImageNet-10, ViT Random Data Forgetting, ImageNet-10, ViT
Retrain $100_{\pm{0.00}}$($0.00$) $100.00_{\pm{0.00}}$($0.00$) $99.97_{\pm{0.03}}$($0.00$) $99.85_{\pm{0.01}}$($0.00$) $\mathbf{0.00}$ $1.41_{\pm{0.06}}$($0.00$) $93.57_{\pm{0.00}}$($0.00$) $99.07_{\pm{0.01}}$($0.00$) $99.27_{\pm{0.01}}$($0.00$) $\mathbf{0.00}$
FT $42.79_{\pm{7.51}}$($57.21$) $40.78_{\pm{10.68}}$($59.22$) $99.96_{\pm{0.01}}$($0.01$) $99.61_{\pm{0.10}}$($0.24$) $\mathbf{29.17}$ $1.38_{\pm{0.16}}$($0.03$) $96.40_{\pm{0.31}}$($2.83$) $99.60_{\pm{0.09}}$($0.53$) $99.10_{\pm{0.30}}$($0.17$) $\mathbf{0.89}$
RL $88.15_{\pm{1.62}}$($11.85$) $96.50_{\pm{1.50}}$($3.50$) $99.93_{\pm{0.02}}$($0.04$) $99.89_{\pm{0.11}}$($0.04$) $\mathbf{3.84}$ $2.62_{\pm{0.80}}$($1.21$) $93.75_{\pm{0.52}}$($0.18$) $98.38_{\pm{0.18}}$($0.69$) $95.05_{\pm{0.26}}$($4.20$) $\mathbf{1.57}$
GA $28.05_{\pm{7.05}}$($71.95$) $58.73_{\pm{5.83}}$($41.27$) $99.97_{\pm{0.01}}$($0.00$) $99.70_{\pm{0.10}}$($0.15$) $\mathbf{28.34}$ $0.82_{\pm{0.16}}$($0.59$) $96.77_{\pm{1.10}}$($3.20$) $99.60_{\pm{0.09}}$($0.53$) $99.53_{\pm{0.09}}$($0.26$) $\mathbf{1.15}$
NegGrad+ $10.90_{\pm{2.87}}$($89.10$) $79.30_{\pm{4.58}}$($20.70$) $99.98_{\pm{0.00}}$($0.01$) $99.78_{\pm{0.00}}$($0.07$) $\mathbf{27.46}$ $1.97_{\pm{0.72}}$($0.56$) $95.48_{\pm{1.67}}$($1.91$) $98.09_{\pm{0.94}}$($0.98$) $98.08_{\pm{0.75}}$($1.19$) $\mathbf{1.16}$
SalUn $93.27_{\pm{1.50}}$($6.73$) $94.00_{\pm{1.00}}$($6.00$) $98.22_{\pm{0.75}}$($1.75$) $98.00_{\pm{0.00}}$($1.85$) $\mathbf{4.08}$ $0.67_{\pm{0.19}}$($0.74$) $95.80_{\pm{0.0}}$($2.23$) $99.65_{\pm{0.07}}$($0.58$) $98.27_{\pm{0.09}}$($0.00$) $\mathbf{1.14}$
SCRUB $99.10_{\pm{0.20}}$($0.90$) $95.67_{\pm{1.77}}$($4.33$) $98.90_{\pm{0.10}}$($1.07$) $99.33_{\pm{0.31}}$($0.52$) $\mathbf{1.71}$ $0.85_{\pm{0.24}}$($0.56$) $95.90_{\pm{0.95}}$($2.33$) $99.58_{\pm{0.23}}$($0.51$) $99.20_{\pm{0.14}}$($0.07$) $\mathbf{0.87}$
Class-F $28.62_{\pm{7.85}}$($71.38$) $55.10_{\pm{2.30}}$($44.90$) $77.50_{\pm{1.62}}$($22.47$) $75.22_{\pm{0.56}}$($24.63$) $\mathbf{40.85}$ n/a n/a n/a n/a n/a
SSD $90.35_{\pm{1.65}}$($9.65$) $60.15_{\pm{1.85}}$($39.85$) $98.43_{\pm{0.55}}$($1.54$) $98.33_{\pm{0.56}}$($1.52$) $\mathbf{13.14}$ $1.12_{\pm{0.04}}$($0.29$) $94.15_{\pm{0.35}}$($0.58$) $98.97_{\pm{0.00}}$($0.10$) $99.40_{\pm{0.00}}$($0.13$) $\mathbf{0.28}$
Ours $95.92_{\pm{0.27}}$($4.08$) $99.40_{\pm{0.14}}$($0.60$) $99.13_{\pm{0.04}}$($0.84$) $99.26_{\pm{0.28}}$($0.59$) $\mathbf{1.53}$ $1.08_{\pm{0.39}}$($0.33$) $91.40_{\pm{1.30}}$($2.17$) $98.97_{\pm{0.35}}$($0.10$) $99.10_{\pm{0.50}}$($0.17$) $\mathbf{0.69}$
Table 1: Performance overview of various MU methods for image classification under two unlearning scenarios on CIFAR-10 using ResNet-18 and ImageNet-10 using ViT. Since Class-F (Kodge et al., 2024) is specifically designed for class-wise forgetting, its results do not apply to random data forgetting scenarios (n/a). Other results are reported in the format $a_{\pm b}$, where $a$ is the mean and $b$ denotes standard deviation $b$ over $10$ independent trials. The performance gap against Retrain is indicated in ($\bullet$), where a lower value is better. $\uparrow$ (or $\downarrow$) indicates that a higher (or lower) value is better. The best performance for each metric is highlighted in green , while the second-best performance is highlighted in red .